40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Aug 2 18:02:46 2010, Jenne, Update, PEM, Seismometer Update: Still not good, but perhpas getting better 
    Reply  Tue Aug 3 01:51:01 2010, Jenne, Update, PEM, Seismometer Update: Still not good, but perhaps getting better 
Message ID: 3349     Entry time: Mon Aug 2 18:02:46 2010     Reply to this: 3351
Author: Jenne 
Type: Update 
Category: PEM 
Subject: Seismometer Update: Still not good, but perhpas getting better 

Today's seismometer diagnosis activities are still underway, this is just an update (since I did some reboots):

Problem 1: X and Z channels on both seismometers were flipped.  I unplugged an X cable (East/West on the cable labels) and the Z channel (vert) would go to floating ADC zero.  Rana suggested that the ADCs sometimes have random channel hopping, and that a reboot of the c0dcu1 computer which handles the PEM ADCU should fix this problem.  I keyed the c0dcu1 / c0daqawg crate, those computers came back just fine, and the channels were no longer flipped.  This is a good thing.  Although now it's actually the Z channels that were / are bad on both seismometers, not the X's.

While rebooting those computers, c1iovme, c1sosvme, c1susvme1 and c1susvme2 crashed.  I rebooted them, although for the first few power cycles c1susvme1&2 couldn't mount /cvs/cds/caltech.  Eventually they did, and life is good again.  Except that the seismometers are still funny.

ELOG V3.1.3-